OpenStudy (anonymous):

I know how to do this, but I don't believe the answer I got. Use Green's Theorem to evaluate the line integral Int(y^3 dx + (x^3 + 3xy^2)dy when C is the path from (0,0) to (1,1) along y = x^3 andfrom (1,1) to (0,0) along y = x. I got 19/20, but it seems like the value should be much lower.

OpenStudy (zarkon):

i get 1/4

OpenStudy (anonymous):

I just redid it and got 1. Can I see your work? I'll post mine in a second.

OpenStudy (zarkon):

\[\int\limits_{0}^{1}\int\limits_{x^3}^{x}\left[(3x^2+3y^2)-(3y^2)\right]dydx\]

OpenStudy (anonymous):

I found my error. Thanks. I'm very rusty. It helps to have someone sanity check me.